OpenStudy (rawfeeds-):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y –3 7 –2 5 Line 2 x y 0 5 2 1 Line 1: y = x + Line 2: y = x +

OpenStudy (rawfeeds-):

@Photon336

rebeccaxhawaii (rebeccaxhawaii):

do you know slope form??

OpenStudy (photon336):

Yeah for this you would have to get the slope first

OpenStudy (photon336):

\[m = (\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} })\] \[(-3,7), (-2,5)\] this means you subtract the y coordinates and the x coordinates separately. \[\frac{ 7-5 }{ ? } = \frac{ 2 }{ ? }\]